# Deciding Subspace reachability problems with application to Skolem’s Problem

### v1

@@ 2.4. Angle evolution, and the Orbit Problem in real projective space
For a matrix $K=\text{Diag}(J_{1},...,J_{N})$ in Jordan canonical form, its $n$th power is given by $J^{n}=\text{Diag}(J_{1}^{n},...,J_{N}^{n})$. Where, for real eigenvalues $\lambda_{i}$ we have
@@
where $D_{i}$ is the dimension of the block $J_{i}$, and (${n\choose k}=0$ if $n<k$. And when $J_{i}$ is a Jordan block associated with a complex conjugate eigenvalue pair expressed in the real canonical form as in Lemma 1, letting
@@
Hence $R$ describes a rotation by the angle $\theta$ followed by multiplication by $|\lambda|$. Letting $H$ denote the $2D_{i}$ dimensional nilpotent matrix of ones just above the principle diagonal, for Jordan block $J_{i}$ associated with a complex conjugate eigenvalue pair, we have
(4) $J_{i}^{n}x=\sum_{i=0}^{D_{i}-2}{n\choose i}|\lambda|^{n-i}\tilde{R}^{n-i}H^{i}x$
@@
By Lemma 3, it suffices to show that $\Gamma(J_{i}^{n}x,P)\rightarrow 1$ as $n\rightarrow\infty$. Recall that $\arccos(\Gamma(J_{i}^{n}x,P))$ is the (minimal) cosine angle between subspaces $J_{i}^{n}x$ and $P$.
Using Lemma 4 we have
@@
where $x=(x_{1},...,x_{\delta})^{\top}$, and the components $x_{i}$ are either $1\times 1$ or $2\times 1$ depending on whether $D$ is $1\times 1$ or $2\times 2$.
Dividing the components of $J_{i}^{n}x$ by $(J_{i}^{n}x)_{1}$ (the first component), and taking a limit we see that
$\lim_{n\rightarrow\infty}\left|\frac{J_{i}^{n}(x)}{(J_{i}^{n}x)_{1}}\right|=(1,0,...,0)^{\top},$
@@
The statement of Lemma 5 did not require $|\lambda_{i}|>1$, because it holds if $|\lambda_{i}|=1$ or $|\lambda_{i}|<1$: when $|\lambda_{i}|=1$, the rate of growth of the first term of the vector is still higher than the others, and when $|\lambda_{i}|<1$, the rate of decline of the first component of the vector is slowest, and hence the statement still holds. Furthermore, the statement continues to hold independent of whether $\lambda_{i}$ is positive or negative — in all cases the angle between the vector and the one or two-dimensional space invariant under $J_{i}$ approaches zero. Finally, we remark that so long as $x$ has at least one non-zero component the lemma holds: after a few iterations all the components of $x$ will be non-zero and the asymptotics kick in accordingly.
@@
Let $J=\text{Diag}(J_{1},...,J_{N})$ be a block diagonal matrix of Jordan blocks, where each $J_{i}$ is associated with a real or complex eigenvalues $\lambda_{i}$, such that $|\lambda_{1}|=|\lambda_{2}|=\cdots=|\lambda_{N}|$. Let $D_{i}$ denote the dimension of $J_{i}$ when $\lambda_{i}$ is real, and half the dimension of $J_{i}$ when $\lambda_{i}$ is complex ($J_{i}$ is the block associated to a complex conjugate pair). Suppose $D_{1}=\cdots=D_{T}>D_{T+1}\geq\cdots\geq D_{N}$, with $T\leq N$.
@@
for any $x$ with at least one non-zero component with respect to every Jordan block composing $J$.
@@
We break this proof into two cases, both of which can be easily combined to obtain the statement. The first case is when $T=1$, and the second case is when $T=N$.
@@
Following the proof of Lemma 5, it is clear that the limit of $|J^{n}x/(J^{n}x)_{1}|$ as $n\rightarrow\infty$ gives $(1,0,...,0)^{\top}$. Thus, all elements $\mathbb{P}x$ where $x$ has at least one non-zero component with respect to $J_{1}$, get taken to $\mathbb{P}P$ in the induced dynamical system in real projective space, where $P$ is either one or two-dimensional depending on whether $J_{1}$ is associated with a real or complex-conjugate pair of eigenvalues, cf. Lemma 5.
@@
We now note that adding many Jordan blocks of the same dimension, all corresponding to the same eigenvalue does not increase the dimension of $P$. This follows from the argument used in proving Lemma 5. Namely, suppose $J$ is a block diagonal matrix composed of $N$ repeated Jordan blocks $J_{i}$, and let $x$ be some vector such that $x_{J_{i}}\neq 0$ for each $J_{i}$. Then take the limit of $|J^{n}x/(J^{n}x)_{j}|$ as $n\rightarrow\infty$, where $|(J^{n}x)_{j}|$ is the largest component of $J^{n}x$. Then the limiting vector will have form $(c_{1},0,...,c_{i},0,...,1,0,...,c_{N},0,...0)^{\top}$, where $0<c_{i}\leq 1$, and the $1$ is in the $j$th position, and the values of the $c_{i}$ can be effectively computed with the coordinates of $x$. Given the above, following the argument of Lemma 5 further, we see that when the repeated Jordan blocks all correspond to a real eigenvalue, then the invariant subspace $P$ orbits approach after projecting to $\mathbb{R}\mathbb{P}^{d-1}$ is of dimension one, and when the repeated Jordan blocks (in real Jordan canonical form) correspond to a complex conjugate pair, $P$ is of dimension two: its elements can easily be written as a linear combination of two vectors.
Hence, in the $T=N$ case, without a loss of generality suppose each Jordan block $J_{i}$ and invariant space $P_{i}$ is paired with a distinct eigenvalue or complex conjugate pair (up to a $-1$ factor), with no Jordan block repeated more than once. Then, as an immediate consequence of the block-diagonal structure of $J$, we have that
@@
where the $P_{i}$ are the one or two-dimensional subspaces invariant under corresponding collections of Jordan blocks associated with the same eigenvalue(s), and $\oplus$ is the direct sum. And, bringing in Lemma 5, we have
@@
Collecting the arguments establishing Cases 1 and 2 above, the statement follows. ∎
@@
In Case 2 of the proof of Lemma 6, we underscore the fact that when there are many Jordan blocks of the same size but that do not correspond to a repeated eigenvalue, but rather the same eigenvalue up to a factor of $-1$, then the orbit $\{\mathbb{P}J^{2n}x\}$ will converge to one subet $\mathbb{P}P$ in projective space, while the orbit $\{\mathbb{P}J^{2n+1}x\}$ will converge to another subset $\mathbb{P}P^{\prime}$ of projective space. This follows from the fact that certain components will change sign with every iteration. However, for the purpose this paper and the proofs of Theorems 1 and 3, we assume non-degeneracy so no two distinct eigenvalues of $A$ will be equal up to a $-1$ factor, and hence this aspect can be safely ignored.
@@
Let $J=\text{Diag}(J_{1},...,J_{N})$ be a block diagonal matrix, where each Jordan block $J_{i}$ is associated with a distinct real eigenvalue or complex conjugate pair. Moreover, suppose $|\lambda_{1}|>|\lambda_{2}|>\cdots>|\lambda_{N}|$. Then there is a subspace $P$ of dimension one (two) whenever $\lambda_{1}$ is real (complex) invariant under $J_{1}$, such that
@@
the $p_{i}(n)$ are polynomials, $l$ is a positive integer, and $n-i=0$ when $i>n$.
Pairing this fact with the statement and proofs of Lemma 5 and Lemma 6, it is clear that $|J^{n}x|/|(J^{n}x)_{1}|\rightarrow(1,0,...0)^{\top}$ as $n\rightarrow\infty$, where $(J^{n}_{x})_{1}$ labels the first component of $J^{n}x$. And thus, orbits under iteration of $J$ induce a sequence in real projective space converging to $\mathbb{P}P$ where $P$ is the one or two dimensional subspace invariant under $J_{1}$, following the arguments of the proofs in Lemmas 5 and 6. ∎
@@ 3. Proofs of Theorems
    By assumption, the target subspace $U$ can be written as
    $U=S_{1}\cup\cdots\cup S_{l}\cup H_{1}\cup\cdots\cdots H_{t},$
    where the $S_{i}$ are subspaces of $\mathbb{Q}^{d}$ represented by sets of rational basis vectors. Similarly, the $H_{t}$ are convex polytopes, and hence are by definition intersections of half-spaces. Then we assume without loss of generality that the $H_{i}$ are presented as collections of hyperplanes (described by rational basis vectors) along with an inequality. We work with this presentation of $U$ as a set of rational bases.
@@
    During the computation of $J$, collect the eigenvalues $\lambda_{1},...,\lambda_{m}$ of $A$, organized so that $|\lambda_{1}|\geq\cdots\geq|\lambda_{m}|$. Suppose $\lambda_{1},...,\lambda_{r}$, $r\leq m$ are the eigenvalues of maximal modulus. Every eigenvalue $\lambda_{i}$ will correspond to one or more Jordan blocks composing $J$, where the number of Jordan blocks associated with an eigenvalue is determined by the algebraic and geometric multiplicity of the eigenvalue, along with the elementary divisors of the characteristic polynomial of $A$.
    Let $P_{A}$ label the minimal polynomial of $A$. Then note that the roots of $P_{A}$ are the eigenvalues of $A$, and the algebraic multiplicity of each root in $P_{A}$ is the dimension of the largest Jordan block associated with the respective root in the Jordan canonical form of $A$. This is a trivial consequence of the definition of the minimal polynomial of $A$. However here we use the real Jordan canonical form, and thus the complex roots of the minimal polynomial will correspond to Jordan blocks of dimension twice their algebraic multiplicity.
    By assumption of Theorem 1, there are $p\leq r$ distinct eigenvalues of maximal modulus with highest algebraic multiplicity in $P_{A}$. As such, by Lemma 6 and the non-degeneracy assumption, there is a single subspace $P$ of dimension $p$ such that $d(\mathbb{P}J^{n}x,\mathbb{P}P)\rightarrow 0$ as $n\rightarrow\infty$ whenever $x$ has non-zero components with respect to the Jordan blocks composing $J$ — which is ensured by the non-triviality assumption placed on $x$ in the theorem statement.
    Return the subspace $P$.
  3. (3)
    The subspace $P$ is taken to be presented as a set of rational basis vectors: it is trivially computable given $x$, along with determining the position of the Jordan blocks of largest dimension corresponding to the eigenvalues of maximal modulus in the matrix $J$. This fact can be readily seen through the proofs of Lemmas 5 and 6.
    Recall $U$ is composed of subspaces $S_{1},...,S_{l}$, and polytopes $H_{1},...,H_{t}$ represented using hyperplanes (subspaces), all taken to be presented with rational bases. Compute $P\cap U$, which can be effectively computed by reducing to computation of the subcases $P\cap S_{i}$ and $P\cap H_{i}$, where $P\cap H_{i}$ is further reduced to computing the intersection of $P$ with the half-spaces composing $H_{i}$. If $P\cap U=\{0\}$, continue. Otherwise halt and terminate execution of the algorithm; this instance of the Orbit Problem cannot be decided by this algorithm.
@@
    $\sigma_{max}=\max\left\{\Gamma(P,Y_{i}):Y_{i}\in\{S_{1},...,S_{l},H_{1},...,H_{t}\}\right\},$
    where the value $\Gamma(P,H_{i})$ corresponds to the cosine of the minimal angle between $P$ and each of the hyperplanes describing a half-space composing $H_{i}$.
@@
    Otherwise, continue until $\Gamma(J^{n}x,P)>\sigma_{max}$: it is an immediate consequence of the proofs of Lemmas 5, 6, 7 that $\Gamma(J^{n}x,P)\rightarrow 1$ monotonically as $n\rightarrow\infty$, and hence by Lemma 3, $d(\mathbb{P}J^{n}x,\mathbb{P}P)\rightarrow 0$ as $n\rightarrow\infty$. But there is an $\epsilon>0$ such that $d(\mathbb{P}P,\mathbb{P}U)=\epsilon$, and hence there is an $N\in\mathbb{N}$ such that for all $n\geq N$, $\mathbb{P}J^{n}x\not\in\mathbb{P}U$. Hence, there is an $N^{\prime}=N^{\prime}(\sigma_{max})\in\mathbb{N}$ such that $J^{n}x\not\in U$ for all $n>N^{\prime}$.
@@
Assume $(A,x,U)$ is a non-degenerate instance of the higher-dimensional Orbit Problem with $A$ nonsingular, and $U$ of dimension one, as in the conditions of the theorem. Immediate from Theorem 1 and its proof, if the orbit $\{A^{n}x\}_{n=0}^{\infty}$ approaches a subspace $W$ after projecting onto real projective space, and $W\cap U=\{0\}$, then the instance is decidable.
Note the condition that $\text{dim}(U)=1$ enforces either $U\subseteq W$ or $U\cap W=\{0\}$. Hence we are left to consider the case when $U\subseteq W$. $A$ is taken to be nonsingular. As a consequence, if $x\not\in W$, $A^{n}x\not\in W$ for all $n$, and hence $A^{n}x\not\in U$ for all $n$ since $W$ is invariant under $A$. If $x\in W$, then consider the reduced, lower-dimensional system $A:W\rightarrow W$ and induct until one of the above conditions is met or $\text{dim}(W)=1$. ∎
@@
Let $\{u_{n}\}_{n=0}^{\infty}$ be an order $d$ non-degenerate linear recurrence sequence with distinct characteristic roots $\lambda_{1},...,\lambda_{m}$, $m\leq d$, ordered so that $|\lambda_{1}|\geq\cdots\geq|\lambda_{m}|$. Suppose the initial terms form a vector $x\in\mathbb{Q}^{d}$ non-trivial with respect to the companion matrix $A$ of the sequence. Then, if $|\lambda_{1}|>|\lambda_{2}|$, it is decidable whether the sequence has a zero term.
@@
Let $A\in\mathbb{Q}^{d\times d}$ label the companion matrix of the LRS $\{u_{n}\}$. Let $x$ denote the non-trivial $d$ dimensional vector of initial terms of the LRS. Then, as noted in Section 2.2, iteration of $A$ over $x$ “shifts” the elements of the LRS through $x$, so that when $x=(u_{d},u_{d-1},...,u_{2},u_{1})^{\top}$, $Ax=(u_{d+1},u_{d},...,u_{3},u_{2})^{\top}$.
Let $E_{1},E_{2},...,E_{d}$ denote the $d$ coordinate subspaces of $\mathbb{R}^{d}$ where the elements of $E_{i}$, $i=1,...,d$, have a 00 in their $i$th component. Then $A^{j}E_{1}=E_{j+1}$, $j=0,...,d-1$. Moreover, $E_{1}\cap E_{2}\cap\cdots\cap E_{d}=\{0\}$. Hence, for any $y\neq 0\in\mathbb{Q}^{d}$, $y\not\in E_{1}\cap\cdots\cap E_{d}$, implying there is an $E_{i}$ such that $y\not\in E_{i}$.
The LRS $\{u_{n}\}$ is taken to have a single dominating real root $\lambda_{1}$. Hence $A$ has a single dominating real root. Taking $A$ to its Jordan canonical form and back via change of basis, we see by consequence of Lemmas 5, 6, 7 that there is a line spanned by some $y\neq 0\in\mathbb{Q}^{d}$, such that $d(\mathbb{P}A^{n}x,\mathbb{P}y)\rightarrow 0$ as $n\rightarrow\infty$. But there is an $E_{i}$ such that $y\not\in E_{i}$, i.e. $\text{span}(y)\cap E_{i}=\{0\}$. Then, appealing to the statement and proof of Theorem 1, it can be decided in a finite number of iterations of $A$ over $x$ whether the LRS has a zero. ∎
Finally, we prove Theorem 3.

### v2

@@ 2.5. Convergence of orbits toward attracting subspaces in $\mathbb{RP}^{d-1}$
For a matrix $K=\text{Diag}(J_{1},\dots,J_{N})$ in Jordan canonical form, its $n$th power is given by $J^{n}=\text{Diag}(J_{1}^{n},\dots,J_{N}^{n})$. Where, for real eigenvalues $\lambda_{i}$ we have
@@
where $D_{i}$ is the dimension of the block $J_{i}$, and ${n\choose k}=0$ if $n<k$. And when $J_{i}$ is a Jordan block associated with a complex conjugate eigenvalue pair expressed in the real canonical form as in Lemma 1, letting
@@
Hence $R$ describes a rotation by the angle $\theta$ followed by multiplication by $|\lambda|$. Letting $H$ denote the $2D_{i}$ dimensional nilpotent matrix of ones just above the principal diagonal, for Jordan block $J_{i}$ associated with a complex conjugate eigenvalue pair, we have
(4) $J_{i}^{n}x=\sum_{i=0}^{D_{i}-1}{n\choose i}|\lambda|^{n-i}\tilde{R}^{n-i}H^{i}x$
@@
By Lemma 3, it suffices to show that $\Gamma(J_{i}^{n}x,P)\rightarrow 1$ as $n\rightarrow\infty$. Recall that $\arccos(\Gamma(J_{i}^{n}x,P))$ is the (minimal) cosine angle between subspaces $J_{i}^{n}x$ and $P$. Using Lemma 4 we have
@@
where $x=(x_{1},\dots,x_{\delta})^{\top}$, and the components $x_{i}$ are either $1\times 1$ or $2\times 1$ depending on whether $D$ is $1\times 1$ or $2\times 2$.
Dividing the components of $J_{i}^{n}x$ by $(J_{i}^{n}x)_{1}$ (the first component), and taking a limit (element-wise) we see that
$\lim_{n\rightarrow\infty}\left|\frac{J_{i}^{n}(x)}{(J_{i}^{n}x)_{1}}\right|=(1,0,\dots,0)^{\top},$
@@
The statement of Lemma 5 did not require $|\lambda_{i}|>1$, because it holds if $|\lambda_{i}|=1$ or $|\lambda_{i}|<1$: when $|\lambda_{i}|=1$, the rate of growth of the first term of the vector is still higher than the others, and when $|\lambda_{i}|<1$, the rate of decline of the first component of the vector is slowest, and hence the statement still holds. Furthermore, the statement continues to hold independent of whether $\lambda_{i}$ is positive or negative — in all cases the angle between the vector and the one or two-dimensional space invariant under $J_{i}$ approaches zero. Finally, we remark that so long as $x$ has at least one non-zero component the lemma holds: after a few iterations all the first component of $J_{i}^{n}x$ will be non-zero and the asymptotics kick in accordingly.
@@
Let $J=\text{Diag}(J_{1},\dots,J_{N})$ be a block diagonal matrix of Jordan blocks, where each $J_{i}$ is associated with a real or complex eigenvalues $\lambda_{i}$, such that $|\lambda_{1}|=|\lambda_{2}|=\cdots=|\lambda_{N}|$. Let $D_{i}$ denote the dimension of $J_{i}$ when $\lambda_{i}$ is real, and half the dimension of $J_{i}$ when $\lambda_{i}$ is complex ($J_{i}$ is the block associated to a complex conjugate pair). Suppose $D_{1}=\cdots=D_{T}>D_{T+1}\geq\cdots\geq D_{N}$, with $T\leq N$.
@@
for any $x$ with at least one non-zero component with respect to the Jordan blocks comprising $J$.
@@
It suffices to prove the statement by analyzing two cases: the first case is when $T=1$, and the second case is when $T=N$. Once these cases have been shown they can be easily combined to obtain the general statement.
@@
Following the proof of Lemma 5, it is clear that the component-wise limit of $|J^{n}x/(J^{n}x)_{1}|$ as $n\rightarrow\infty$ gives $(1,0,\dots,0)^{\top}$. Thus, all elements $\mathbb{P}x$ where $x$ has at least one non-zero component with respect to $J_{1}$, get taken to $\mathbb{P}P$ in the induced dynamical system in real projective space, where $P$ is either one or two-dimensional depending on whether $J_{1}$ is associated with a real or complex-conjugate pair of eigenvalues, cf. Lemma 5.
@@
We first note that adding many Jordan blocks of the same dimension, all corresponding to the same eigenvalue does not increase the dimension of $P=P(x,J)$. This follows from the argument used in proving Lemma 5. Namely, suppose $J$ is a block diagonal matrix composed of $N$ repeated Jordan blocks $J_{i}$, and let $x$ be some vector such that $x_{J_{i}}\neq 0$ for each $J_{i}$. Then take the limit of $|J^{n}x/(J^{n}x)_{j}|$ as $n\rightarrow\infty$, where $|(J^{n}x)_{j}|$ is the largest component of $J^{n}x$. Then the limiting vector will have form $(c_{1},0,\dots,c_{i},0,\dots,1,0,\dots,c_{N},0,\dots 0)^{\top}$, where $0\leq c_{i}\leq 1$, and the $1$ is in the $j$th position, and the values of the $c_{i}$ can be effectively computed with the coordinates of $x$ through taking appropriate ratios of the components of $x$. Hence, when the repeated Jordan blocks all correspond to a real eigenvalue, the invariant subspace $P$ is of dimension one. And when the repeated Jordan blocks correspond to a complex conjugate pair, $P$ is of dimension two.
Hence provided with this fact, in the $T=N$ case we may assume without any loss of generality that each Jordan block $J_{i}$ (and therefore invariant space $P_{i}$) is paired with a distinct eigenvalue or complex conjugate pair (up to a 1-1 factor), with no Jordan block repeated more than once. Then, as an immediate consequence of the block-diagonal structure of $J$, we have that
@@
where the $P_{i}$ are the one or two-dimensional subspaces invariant under corresponding collections of Jordan blocks associated with the same eigenvalue(s), and $\oplus$ is the direct sum. Using Lemma 5, we have
@@
Collecting the arguments establishing Cases 1 and 2 above, the statement follows: the subspace the dynamics are asymptotic to is determined by the Jordan blocks of largest dimension. ∎
@@
In Case 2 of the proof of Lemma 6, we emphasize that as long as $x$ has one non-zero component with respect to each Jordan block then the proof goes through. What our proof perhaps leaves unclear in regard to this matter is that in the limit discussed in the second paragraph, we have $0\leq c_{i}\leq 1$, and hence it is possible that a $c_{i}=0$. But this is fully allowed, since in this case the particular orbit will still approach $P$, even though it is simply approaching a particular subset of $P$.
Remark 3.
We also note that in Case 2 of the proof of Lemma 6, we underscore the fact that when there are many Jordan blocks of the same size but that do not correspond to a repeated eigenvalue, but rather the same eigenvalue up to a factor of 1-1, then the orbit $\{\mathbb{P}J^{2n}x\}$ will converge to one subset $\mathbb{P}P$ in projective space, while the orbit $\{\mathbb{P}J^{2n+1}x\}$ will converge to another subset $\mathbb{P}P^{\prime}$ of projective space. This follows from the fact that certain components will change sign with every iteration. However, for the purpose this paper and the proofs of Theorems 1 and 3, we assume non-degeneracy so no two distinct eigenvalues of $A$ will be equal up to a 1-1 factor, and hence this aspect can be safely ignored.
@@
Let $J=\text{Diag}(J_{1},\dots,J_{N})$ be a block diagonal matrix, where each Jordan block $J_{i}$ is associated with a distinct real eigenvalue or complex conjugate pair. Moreover, suppose $|\lambda_{1}|>|\lambda_{2}|>\cdots>|\lambda_{N}|$. Then there is a subspace $P$ of dimension one (two) whenever $\lambda_{1}$ is real (complex) invariant under $J_{1}$, such that
@@
the $p_{i}(n)$ are polynomials, $l$ is a positive integer, and ni=0n-i=0 when $i>n$.
Pairing this fact with the statement and proofs of Lemma 5 and Lemma 6, it is clear that $|J^{n}x|/|(J^{n}x)_{1}|\rightarrow(1,0,\dots 0)^{\top}$ as $n\rightarrow\infty$, where $(J^{n}_{x})_{1}$ labels the first component of $J^{n}x$. And thus, orbits under iteration of $J$ induce a sequence in real projective space converging to $\mathbb{P}P$ where $P$ is the one or two dimensional subspace invariant under $J_{1}$, following the arguments of the proofs in Lemmas 5 and 6. ∎
@@ 3. Proofs of Theorems
    During the computation of $J$, collect the eigenvalues $\lambda_{1},\dots,\lambda_{m}$ of $A$, organized so that $|\lambda_{1}|\geq\cdots\geq|\lambda_{m}|$. Suppose $\lambda_{1},\dots,\lambda_{r}$, $r\leq m$ are the eigenvalues of maximal modulus. Every eigenvalue $\lambda_{i}$ will correspond to one or more Jordan blocks composing $J$, where the number of Jordan blocks associated with an eigenvalue is determined by the algebraic and geometric multiplicity of the eigenvalue, along with the elementary divisors of the characteristic polynomial of $A$.
    Let $P_{A}$ label the minimal polynomial of $A$. Then note that the roots of $P_{A}$ are the eigenvalues of $A$, and the multiplicity of each root in $P_{A}$ is the dimension of the largest Jordan block associated with the respective root in the Jordan canonical form of $A$. This is a trivial consequence of the definition of the minimal polynomial of $A$. However here we use the real Jordan canonical form, and thus the complex roots of the minimal polynomial will correspond to Jordan blocks of dimension twice their multiplicity.
    By assumption of Theorem 1, there are $p\leq r$ distinct eigenvalues of maximal modulus with highest multiplicity in $P_{A}$. As such, by Lemma 6 and the non-degeneracy assumption, there is a single subspace $P$ of dimension $p$ such that $d(\mathbb{P}J^{n}x,\mathbb{P}P)\rightarrow 0$ as $n\rightarrow\infty$ whenever $x$ has non-zero components with respect to the Jordan blocks composing $J$ — which is ensured by the non-triviality assumption placed on $x$ in the theorem statement.
    Return the subspace $P=P(J,x)$.
@@
    The subspace $P$ is taken to be presented as a set of rational basis vectors: it is computable given $x$, along with determining the position of the Jordan blocks of largest dimension corresponding to the eigenvalues of maximal modulus in the matrix $J$. This fact follows from the proofs of Lemmas 5 and 6; that is, $P$ is simply a direct sum of either standard basis vectors in the Jordan basis, or invariant subspaces described by rational linear combinations of standard basis vectors, the coefficients of which are effectively computable from the entries of $x$ as commented in Lemma 6.
    Compute $P\cap U$. If $P\cap U=\{0\}$, continue. Otherwise halt and terminate execution of the algorithm; this instance of the Orbit Problem cannot be decided by this algorithm.
@@
    $\sigma_{max}=\Gamma(P,U).$
    where the value $\Gamma(P,U)$ corresponds to the cosine of the minimal angle between $P$ and $U$.
@@
    Otherwise, continue until $\Gamma(J^{n}x,P)>\sigma_{max}$: since $\Gamma(J^{n}x,P)\rightarrow 1$ as $n\rightarrow\infty$, by Lemma 3 $d(\mathbb{P}J^{n}x,\mathbb{P}P)\rightarrow 0$ as $n\rightarrow\infty$. But there is an $\epsilon>0$ such that $d(\mathbb{P}P,\mathbb{P}U)=\epsilon$, and hence there is an $N\in\mathbb{N}$ such that for all $n\geq N$, $\mathbb{P}J^{n}x\not\in\mathbb{P}U$. Hence, there is an $N^{\prime}=N^{\prime}(\sigma_{max})\in\mathbb{N}$ such that $J^{n}x\not\in U$ for all $n>N^{\prime}$.
@@
As usual, if the instance is degenerate we can reduce to a finite number of non-degenerate instances to check. Hence we assume non-degeneracy. We begin by arguing the case when $A$ is nonsingular.
Take $(A,x,U)$ is a non-degenerate instance of the higher-dimensional Orbit Problem with $A$ nonsingular and $U$ of dimension one. In addition, recall that it is assumed that the modulus of the eigenvalues of $A$ are all distinct, up to complex conjugate pairs. Note that immediate from Theorem 1 and its proof, if the orbit $\{A^{n}x\}_{n=0}^{\infty}$ approaches a subspace $W$ after projecting onto real projective space, and $W\cap U=\{0\}$, then the instance is decidable. Moreover, the condition that $\text{dim}(U)=1$ enforces either $U\subseteq W$ or $U\cap W=\{0\}$. Hence, we are left to consider the case when $U\subseteq W$.
AA is taken to be nonsingular with eigenvalues of distinct modulus up to conjugate pairs, so the dimension of the attracting subspace $W$ is strictly less than that of the ambient space, unless $d=2$ with $A$ admitting a complex conjugate eigenvalue pair. But in such a situation $A$ merely acts as a rational or irrational rotation of the plane up to scaling, and with $U$ one-dimensional this problem is well known to be decidable.
Suppose $x\not\in W$: then since $A$ is nonsingular $A^{n}x\not\in U$ for all $n$ since $W$ is invariant under $A$. If $x\in W$, then consider the reduced, lower-dimensional system $A:W\rightarrow W$ and induct until one of the above conditions is met or $\text{dim}(W)=1$ (halt and return “yes”) or $\text{dim}(W)=2$ (consider the circle rotation). Notice we can continue this induction since the eigenvalues of $A$ have distinct modulus.
Finally, note that if $A$ is singular, decidability is proved via essentially identical argument. Namely, in the singular case $A$ collapses points to some subspace $Q$, which is readily computed. Check if $x$ or $Ax\in U$. If not, via basic argument one may consider the restricted system $A|_{Q}$ with $U\cap Q$, and ask the orbit problem in this lower-dimensional setting, then proceed as before. ∎
@@
Let $\{u_{n}\}_{n=0}^{\infty}$ be an order $d$ non-degenerate linear recurrence sequence with distinct characteristic roots $\lambda_{1},\dots,\lambda_{m}$, $m\leq d$, ordered so that $|\lambda_{1}|\geq\cdots\geq|\lambda_{m}|$. Suppose the initial terms form a vector $x\in\mathbb{Q}^{d}$ non-trivial with respect to the companion matrix $A$ of the sequence. Then, if $|\lambda_{1}|>|\lambda_{2}|$, it is decidable whether the sequence has a zero term.
@@
Let $A\in\mathbb{Q}^{d\times d}$ label the companion matrix of the LRS $\{u_{n}\}$. Let $x$ denote the non-trivial $d$ dimensional vector of initial terms of the LRS. Then, as noted in Section 2.2, iteration of $A$ over $x$ “shifts” the elements of the LRS through $x$, so that when $x=(u_{d},u_{d-1},\dots,u_{2},u_{1})^{\top}$, $Ax=(u_{d+1},u_{d},\dots,u_{3},u_{2})^{\top}$.
Let $E_{1},E_{2},\dots,E_{d}$ denote the $d$ coordinate subspaces of $\mathbb{R}^{d}$ where the elements of $E_{i}$, $i=1,\dots,d$, have a 0 in their $i$th component. Then $AE_{d}\subseteq E_{d-1},AE_{d-1}\subseteq E_{d-2},\dots,AE_{2}\subseteq E_{1}$. Moreover, $E_{1}\cap E_{2}\cap\cdots\cap E_{d}=\{0\}$. Hence, for any $y\neq 0\in\mathbb{Q}^{d}$, $y\not\in E_{1}\cap\cdots\cap E_{d}$, implying there is an $E_{i}$ such that $y\not\in E_{i}$.
The LRS $\{u_{n}\}$ is taken to have a single dominating real root $\lambda_{1}$. Hence $A$ has a single dominating real root. Taking $A$ to its Jordan canonical form and back via a change of basis, we see by consequence of Lemmas 5, 6, 7 that there is a line spanned by some $y\neq 0\in\mathbb{Q}^{d}$, such that $d(\mathbb{P}A^{n}x,\mathbb{P}y)\rightarrow 0$ as $n\rightarrow\infty$. But there is an $E_{i}$ such that $y\not\in E_{i}$, i.e. $\text{span}(y)\cap E_{i}=\{0\}$. Then, appealing to the statement and proof of Theorem 1, it can be decided in a finite number of iterations of $A$ over $x$ whether the LRS has a zero. ∎
We now prove Theorem 3.